What is the probability that no more than 156 U.S adults of the 240 surveyed believe solar energy is very effective in minimizing air pollution?
1
Expert's answer
2020-04-27T19:53:15-0400

"Let\;p=\frac{156}{240}=0.65\\
np=0.65\times 240=156, 
\\n(1-p)=0.35\times 240=84,\\
np\geq 5, n(1-p)\geq5, \\
\text{so we can use the Normal Distribution}\\
\text{ to Approximate Binomial Probabilities} 
\\p(x \leq 156)=p(z \leq \frac {156.5-240\times 0.65}{\sqrt{240(0.65)(0.35)}})\\
p(z \leq 0.07)=0.5+p(0\leq z\leq 0.07)\\
=0.5+0.0279=0.5279"